### v1

# Optimistic Safety for Linearly-Constrained Online Convex Optimization

Abstract
The setting of online convex optimization (OCO) under unknown constraints has garnered significant attention in recent years. In this work, we consider a version of this problem with static linear constraints that the player receives noisy feedback of and must always satisfy. By leveraging our novel design paradigm of optimistic safety, we give an algorithm for this problem that enjoys $\tilde{\mathcal{O}}(\sqrt{T})$ regret. This improves on the previous best regret bound of $\tilde{\mathcal{O}}(T^{2/3})$ while using only slightly stronger assumptions of independent noise and an oblivious adversary. Then, by recasting this problem as OCO under time-varying stochastic linear constraints, we show that our algorithm enjoys the same regret guarantees in such a setting and never violates the constraints in expectation. This contributes to the literature on OCO under time-varying stochastic constraints, where the state-of-the-art algorithms enjoy $\tilde{\mathcal{O}}(\sqrt{T})$ regret and $\tilde{\mathcal{O}}(\sqrt{T})$ violation when the constraints are convex and the player receives full feedback. Additionally, we provide a version of our algorithm that is more computationally efficient and give numerical experiments comparing it with benchmark algorithms.
1 Introduction
Online convex optimization (OCO), formalized by Zinkevich (2003), is a setting where, in each round, a player chooses a vector action, suffers the cost of that action according to an adversarily-chosen cost function, and subsequently observes the cost function. This setting has found broad applicability in various fields including online advertising McMahan et al. (2013), network routing Awerbuch & Kleinberg (2008), portfolio management Cover (1991) and resource allocation Yu & Neely (2019). However, despite its significance, this fundamental setting assumes that any constraints on the player’s actions are known a priori, which is not the case in many real-world applications. As a result, the problem of OCO under unknown constraints has received considerable attention in the last decade, e.g. Mannor et al. (2009); Yu et al. (2017); Cao & Liu (2018); Guo et al. (2022).
Two distinct directions have emerged in the literature on OCO under unknown constraints. The first direction considers time-varying constraints that are revealed after the player commits to an action, while the second direction considers unknown static constraints that the player receives incomplete feedback on. When the constraints are time-varying, the player cannot hope to always satisfy the constraint so it is instead natural to bound the cumulative constraint violation [^1] Here, we use $g_{t}$ to refer to the constraint in the $t$th round. $\sum_{t=1}^{T}g_{t}(x_{t})$ that the player incurs. However, it has been shown that sublinear regret and sublinear violation cannot be simultaneously achieved Mannor et al. (2009), so several works have instead focused on the case with iid constraints and provided algorithms with $\tilde{\mathcal{O}}(\sqrt{T})$ regret and $\tilde{\mathcal{O}}(\sqrt{T})$ violation Yu et al. (2017); Wei et al. (2020); Castiglioni et al. (2022). On the other hand, the setting with OCO under fixed unknown constraints enjoys $\tilde{\mathcal{O}}(T^{2/3})$ regret and zero instantaneous constraint violation (i.e. $g(x_{t})\leq 0\ \forall t$) when the constraints are linear and the player receives noisy constraint feedback at the chosen actions Chaudhary & Kalathil (2022); Chang et al. (2023).
Ref. Const. Var. Const. Type Const. Feedback Regret Cumul. Viol. Instant. Viol.
[1],[2] static linear noisy bandit $\tilde{\mathcal{O}}(T^{2/3})$ 00
Thm. 4.1 static linear noisy bandit $\tilde{\mathcal{O}}(\sqrt{T})$ 00
[3],[4],[5] stochastic convex full $\tilde{\mathcal{O}}(\sqrt{T})$ $\tilde{\mathcal{O}}(\sqrt{T})$
Prop. 5.1 stochastic linear bandit $\tilde{\mathcal{O}}(\sqrt{T})$ 00
Table 1: Existing work and novel contributions on OCO under unknown static constraints and OCO under stochastic constraints. Our contributions are highlighted in gray. Prior work is referenced as: [1] = Chaudhary & Kalathil (2022), [2] = Chang et al. (2023), [3] = Yu et al. (2017), [4] = Wei et al. (2020), [5] = Castiglioni et al. (2022). In the “instantaneous violation” column, the stated guarantees hold with high probability for static constraints and in expectation for stochastic constraints.
We contribute to this literature as summarized in Table 1 and in the following:
  • We give the first algorithm with $\tilde{\mathcal{O}}(\sqrt{T})$ regret for OCO under unknown linear constraints with noisy feedback. With only slightly stronger assumptions of independent noise and an oblivious adversary, this improves on the best existing $\tilde{\mathcal{O}}(T^{2/3})$ regret attained by Chaudhary & Kalathil (2022) and Chang et al. (2023).
  • We show that, with slight modification, our algorithm enjoys $\tilde{\mathcal{O}}(\sqrt{T})$ regret and zero instantaneous violation under stochastic linear constraints and bandit constraint feedback. To put this in context, the state-of-the-art guarantees for OCO under stochastic constraints are $\tilde{\mathcal{O}}(\sqrt{T})$ regret and $\tilde{\mathcal{O}}(\sqrt{T})$ cumulative violation Yu et al. (2017); Castiglioni et al. (2022) when the constraints are convex and the player recieves full feedback.
These contributions are thanks to our novel approach of using optimistic estimates of the constraints (outer-approximations of the constraint set) to identify actions that are low-regret, but potentially unsafe, and then using pessimistic estimates of the constraints (inner-approximations of the constraint set) to find the scalings on these actions that will ensure constraint satisfaction. Ultimately, this results in decision-making that is simultaneously low-regret and safe. We call this design paradigm optimistic safety and foresee it being broadly applicable to learning and optimization under uncertain constraints.
1.1 Overview
In Section 2, we define relevant notation and then give the problem setup and assumptions. Then, we design a meta-algorithm in Section 3 that leverages the idea of optimistic safety by using a generic online optimization algorithm to choose actions from the optimistic action set and then scales these actions down into the pessimistic set before playing them. In Section 4, we show that this meta-algorithm ensures $\tilde{\mathcal{O}}(d\sqrt{T})$ regret and zero constraint violation. We then recast the problem as OCO under stochastic constraints in Section 5, and show that our proposed approach enjoys $\tilde{\mathcal{O}}(d\sqrt{T})$ regret guarantees and zero expected constraint violation. In Section 6, we address computational inefficiencies in our original algorithm by giving a version that only requires $2d$ convex projections in each round and enjoys $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret. Lastly, we give simulation results in Section 7 that verify the theoretical claims and show improved performance compared to baseline approaches.
1.2 Related work
This work is related to works on OCO with unknown linear constraints, OCO with time-varying constraints, projection-free OCO, safe learning and online control. We provide separate discussion for each of these fields in the following.
OCO with Unknown Constraints
The problem of OCO under unknown linear constraints and noisy feedback was first studied by Chaudhary & Kalathil (2022) who gave an algorithm with $\tilde{\mathcal{O}}(T^{2/3})$ regret guarantees. In comparison to Chaudhary & Kalathil (2022), we use the slightly stronger assumptions of independent noise and oblivious adversary as discussed in Section 2.3. The approach taken by Chaudhary & Kalathil (2022) was then extended to the distributed and nonconvex setting in Chang et al. (2023), where an $\tilde{\mathcal{O}}(T^{2/3}+T^{1/3}C_{T})$ regret bound was given with $C_{T}$ as the path length of the best minimizer sequence.
OCO with Time-varying Constraints
For OCO with time-varying constraints, Mannor et al. (2009) showed that no algorithm can simultaneously achieve both sublinear constraint violation and sublinear regret against the best action that satisfies the constraints on average. Following this negative result, considerable effort has been focused on ways to weaken the setting to guarantee sublinear regret and violation, e.g. Neely & Yu (2017); Sun et al. (2017); Liakopoulos et al. (2019); Cao & Liu (2018); Yi et al. (2022); Guo et al. (2022); Castiglioni et al. (2022). Most relevantly, several works study the setting in which the constraints are bounded convex functions that are sampled iid and guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ regret and cumulative violation Yu et al. (2017); Wei et al. (2020); Castiglioni et al. (2022). In Section 5, we guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ regret and zero expected instantaneous violation in a related setting where the constraints are linear with subgaussian tails and the player receives bandit feedback of the constraint. Note that our assumptions of subgaussianity and bandit feedback are weaker than bounded and full feedback as considered by Yu et al. (2017), Wei et al. (2020) and Castiglioni et al. (2022), although our assumption of linearity is stronger than their assumption of convexity.
Projection-free OCO
The problem of OCO under unknown constraints has similar goals to that of projection-free OCO, which is focused on methods for OCO that do not require the costly projection operation required for classical methods, such as OGD Zinkevich (2003). A large number of projection-free OCO methods use the cheaper linear optimization oracle instead of the projection oracle and often use variants of the Franke-Wolfe method, e.g. Hazan & Kale (2012); Garber & Hazan (2016); Hazan & Minasyan (2020); Kretzu & Garber (2021). Another line of work avoids projections by allowing some constraint violation and uses similar approaches to OCO with time-varying constraints, e.g. Mahdavi et al. (2012); Yu & Neely (2020); Guo et al. (2022). Some other approaches use the membership oracle Mhammedi (2022) or the constraint function and gradient at any point Levy & Krause (2019). The literature on projection-free OCO is distinct from OCO under unknown constraints in that they either allow constraint violation or consider different types of oracles that use the constraints, e.g. membership oracle, constraint function value and gradient at any point, or linear optimization oracle.
Safe Learning
Our setting falls under the umbrella of “safe learning,” which often deals with learning problems that have uncertain constraints. Some examples of such problems are constrained MDPs Achiam et al. (2017); Wachi & Sui (2020); Liu et al. (2021); Amani et al. (2021), safe Gaussian process bandits Sui et al. (2015, 2018), safe optimization with uncertain constraints Usmanova et al. (2019); Fereydounian et al. (2020) and safe linear bandits Amani et al. (2019); Moradipari et al. (2021); Pacchiano et al. (2021). The most relevant to our setting is safe linear bandits Moradipari et al. (2021); Pacchiano et al. (2021) as the two settings have the same constraint feedback and both concern regret minimization. However, the safe linear bandit is distinct from the setting we consider as it has stochastic linear cost functions and bandit feedback versus the adversarial convex cost functions and full cost feedback used in our setting. Also, note that the setting we study in the present work generalizes the safe optimization setting in Usmanova et al. (2019) as they consider the same constraints and constraint feedback with a static cost function.
Online Control
There is an interesting connection between OCO with unknown linear constraints and online control with unknown linear dynamics (e.g. Cohen et al. (2018); Simchowitz et al. (2020); Cassel et al. (2022)). In particular, both settings have unknown linear functions with noisy feedback and adversarial convex costs. However, in online control, the unknown linear function impacts the incurred cost, while in OCO with unknown linear constraints, the unknown linear function forms the constraint. The latter setting poses a fundamentally different challenge that we aim to address in this work.
2 Preliminaries
2.1 Notation
[…]
2.2 Problem Setup
We study an online convex optimization problem with an unknown linear constraint. This problem is played over $T$ rounds, where in each round $t\in[T]$, a player chooses an action $x_{t}$ from the convex action set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and, subsequently, suffers the cost $f_{t}(x_{t})$ according to an adversarily-chosen convex function $f_{t}:\mathcal{X}\rightarrow\mathbb{R}$. The player then observes the cost function $f_{t}$ and receives noisy feedback on the constraint $Ax_{t}+\epsilon_{t}$ where $A\in\mathbb{R}^{n\times d}$ is the constraint matrix and $\epsilon_{t}$ is a noise term.
The player aims to ensure that the constraint $Ax_{t}\leq b$ is always satisfied despite the fact that the constraint matrix $A$ is unknown. At the same time, the player also aims to minimize her regret with respect to the best constraint-satisfying action in hindsight, i.e.
$R_{T}:=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star}),$
where $x^{\star}\in\operatorname*{arg\,min}_{x\in\mathcal{Y}}\sum_{t=1}^{T}f_{t}(x)$ and $\mathcal{Y}=\{x\in\mathcal{X}:Ax\leq b\}$.
2.3 Assumptions
Our approach to this setting relies on the following assumptions. We first assume that the cost functions have bounded gradients and the action set is compact, which are standard in OCO (e.g. Zinkevich (2003); Hazan et al. (2016))
Assumption 2.1 (Bounded gradients).
For all $t\in[T]$, $f_{t}$ is differentiable and $\|\nabla f_{t}(x)\|\leq G$ for all $x\in\mathcal{X}$.
Assumption 2.2 (Compact action set).
The action set $\mathcal{X}$ is closed and $\|x\|\leq D/2$ for all $x\in\mathcal{X}$.
We then assume that the noise terms are independent and subgaussian, the constraint matrix is bounded and that there is an initial safe action that is known. These match the assumptions used in Chaudhary & Kalathil (2022); Chang et al. (2023), except that we use the mild additional assumption that the noise terms are independent.
Assumption 2.3 (Subgaussian noise).
The noise $(\epsilon_{t})_{t\in[T]}$ is a sequence of independent random variables that are element-wise $\rho$-subgaussian, i.e. it holds for all $t\in[T]$ and $i\in[n]$ that $\mathbb{E}[\epsilon_{t,i}]=0$ and $\mathbb{E}[\exp(\lambda\epsilon_{t,i})]\leq\exp(\frac{\lambda^{2}\rho^{2}}{2}),\forall\lambda\in\mathbb{R}$ where $\epsilon_{t,i}$ is the $i$th element of $\epsilon_{t}$.
Assumption 2.4 (Bounded constraint).
Each row of $A$ is bounded, i.e. $\|a_{i}\|\leq S$ for all $i\in[n]$ where $a_{i}^{\top}$ is the $i$th row of $A$.
Assumption 2.5 (Initial safe action).
[…]
Lastly, we assume that the adversary is oblivious such that the cost functions are chosen independently of the player’s actions and the noise terms.
Assumption 2.6 (Oblivious adversary).
The adversary is oblivious and therefore the cost functions are fixed in advance.
Algorithm 1 Optimistically Safe OCO (OSOCO)
0:  Action set $\mathcal{X}$, Horizon $T$, Online Algorithm $\mathcal{A}$, Confidence set radius $\beta_{t}$, Regularization $\lambda$.
1:  Initialize: $t=1,j=1,V_{t}=\lambda I,S_{t}=\mathbf{0}$.
2:  while $t\leq T$ do
3:     Initialize phase: $\bar{\beta}_{j}=\beta_{t},\bar{S}_{j}=S_{t},\bar{V}_{j}=V_{t}$.
4:     Estimate constraints: $\hat{A}_{j}=\bar{S}_{j}\bar{V}_{j}^{-1}$.
5:     Update pessimistic action set: $\mathcal{Y}_{j}^{p}=\left\{x\in\mathcal{X}:\hat{A}_{j}x+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq b\right\}$
6:     Update optimistic action set: $\mathcal{Y}_{j}^{o}:=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq b\right\}$
7:     Initialize $\mathcal{A}$ with $\mathcal{Y}_{j}^{o}$ as action set.
8:     while $\det(V_{t})\leq 2\det(\bar{V}_{j})$ and $t\leq T$  do
9:        Receive $\tilde{x}_{t}$ from $\mathcal{A}$.
10:        Find safe scaling: $\gamma_{t}=\max\left\{\mu\in\left[0,1\right]:\mu\tilde{x}_{t}\in\mathcal{Y}_{j}^{p}\right\}$.
11:        Play $x_{t}=\gamma_{t}\tilde{x}_{t}$ and observe $f_{t},y_{t}$.
12:        Send $f_{t}$ to $\mathcal{A}$.
13:        $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$.
14:        $S_{t+1}=S_{t}+y_{t}x_{t}^{\top}$.
15:        $t=t+1$.
16:     end while
17:     Terminate $\mathcal{A}$.
18:     $j=j+1$.
19:  end while
3 Algorithm
To address the stated problem, we propose the algorithm Optimistically Safe OCO (OSOCO) given in Algorithm 1. The key idea behind this algorithm is that it maintains both an optimistic action set, which overestimates the feasible set, and a pessimistic action set, which underestimates the feasible set, to simultaneously ensure low regret and constraint satisfaction. Since the optimistic set contains the true feasible set and therefore the optimal action, we ensure low regret by running an online optimization algorithm over the optimistic set. We then scale down the action selected by the online optimization algorithm in each round until it is in the pessimistic set and therefore satisfies the constraints. Notably, our approach of performing regret minimization while simultaneously learning the constraints differs from existing approaches (e.g. Chaudhary & Kalathil (2022); Chang et al. (2023)) which first perform pure exploration to learn the constraints and then perform regret minimization.
We give the details on OSOCO in Section 3.1. Then, in Section 3.2, we give the requirements on the online optimization algorithm $\mathcal{A}$ that is instantiated by OSOCO.
3.1 Details on OSOCO
The OSOCO algorithm operates over a number of phases, indexed by $j$. We describe the initialization of each phase, the update in each round, and then the phase termination criteria in the following sections.
Phase Initialization (lines 3-7)
At the beginning of each phase $j$, the regularized least-squares estimator of the constraint $\hat{A}_{j}$ is calculated (line 4), which is then used to construct the pessimistic action set $\mathcal{Y}_{j}^{p}$ (line 5) and the optimistic action set $\mathcal{Y}_{j}^{o}$ (line 6). Then, the online optimization algorithm $\mathcal{A}$ is initialized with the optimistic action set (line 7). To appropriately construct the pessimistic and optimistic action sets, we use the confidence set for least-squares estimators from Abbasi-Yadkori et al. (2011), which is given in the following theorem.
Theorem 3.1 (Theorem 2 in Abbasi-Yadkori et al. (2011)).
Fix some $\delta\in(0,1)$ and let $\check{A}_{t}=S_{t}V_{t}^{-1}$. Then, with probability at least $1-\delta$, it holds for all $t\geq 1$ and $x\in\mathbb{R}^{d}$ that
$(\check{A}_{t}-A)x\in\beta_{t}\|x\|_{V_{t}^{-1}}\mathbb{B}_{\infty}$
where
$\beta_{t}:=\rho\sqrt{d\log\left(\frac{1+(t-1)D^{2}/\lambda}{\delta/n}\right)}.$ (1)
It immediately follows from the validity of this confidence set that the optimistic set (resp. pessimistic set) is a superset (resp. subset) of the true feasible set for all rounds with high probability, i.e. $\mathcal{Y}_{j}^{p}\subseteq\mathcal{Y}\subseteq\mathcal{Y}_{j}^{o}$ for all $j$ with probability at least $1-\delta$.
Update in Each Round (lines 9-14)
In each round, the online optimization algorithm $\mathcal{A}$ chooses a point $\tilde{x}_{t}$ from the optimistic set (line 9), and then the OSOCO algorithm finds the largest scaling $\gamma_{t}$ on $\tilde{x}_{t}$ such that it is in the pessimistic set (line 10), before playing $x_{t}=\gamma_{t}\tilde{x}_{t}$ (line 11). The matrices used for estimation ($V_{t},S_{t}$) are then updated (lines 13). The reason that we scale the optimistic action $\tilde{x}_{t}$ down into the pessimistic set rather than using another method, such as the euclidean projection, is because scaling this action down ensures that we are still choosing an action in the direction of the optimistic action which gives the algorithm more constraint information in this direction. This ultimately leads to effective decision-making because the optimistic actions are updated on the optimistic set and therefore will naturally gravitate towards areas that are uncertain and/or areas that have low regret, effectively balancing exploitation and exploration.
Phase Termination
The phase ends when the determinant of the Gram matrix ($V_{t+1}$) has doubled (line 8), after which it terminates $\mathcal{A}$ (line 17) and starts a new phase. This type of “rare update” ensures that the phase only ends when a significant amount of new constraint information has been collected. As such, it results in a small number $\mathcal{O}(d\log(T))$ of phases and therefore there will only be a small cost to restarting the OCO algorithm in each phase. Note that a related form of rare updating was used as early as Abbasi-Yadkori et al. (2011) to reduce the number of determinant calculations in stochastic linear bandits.
3.2 Requirements of Online Algorithm $\mathcal{A}$
The online optimization algorithm $\mathcal{A}$ needs to be chosen from a specific class of algorithms in order for OSOCO to work. In particular, the optimistic action set is not necessarily convex and $\mathcal{A}$ can be terminated at any time, so $\mathcal{A}$ needs to perform well when its action set is nonconvex and when the horizon is unknown. We make this concrete in the following.
Definition 3.2 (Regret Bound on $\mathcal{A}$).
Consider an online optimization setting where in each round $t$, the player chooses $x_{t}$ from the action set $\mathcal{X}$, suffers the cost $f_{t}(x_{t})$ and then observes $f_{t}$. Importantly,
  • $\mathcal{X}$ is not necessarily convex,
  • the game can be terminated at any time (and therefore the final round $T$ is not known to the player).
Also, assume that all $f_{t}$ are convex and that Assumptions 2.1, 2.2 and 2.6 hold. We further assume that $\mathcal{A}$ chooses a distribution $p_{t}$ in each round, which it uses to sample $x_{t}$. [^3] We assume that $\mathcal{A}$ is randomized because an online optimization algorithm needs randomness to achieve sublinear regret when the action set is not convex, see Cesa-Bianchi & Lugosi (2006). Then, $C_{\mathcal{A}}$ is defined as a quantity such that
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x_{t})\leq C_{\mathcal{A}}\sqrt{T},$
if the algorithm $\mathcal{A}$ is played in any such online setting. We let $C_{\mathcal{A}}$ depend on $T$ at most $\mathcal{O}(\mathrm{polylog}(T))$.
Most algorithms that admit such a regret bound are those designed for the online nonconvex optimization setting as they can handle the nonconvex action set (and then the “doubling trick” can be used to handle the unknown horizon). Some examples of applicable algorithms are FTPL Kalai & Vempala (2005); Agarwal et al. (2019); Suggala & Netrapalli (2020) and continuous versions of the Hedge algorithm Freund & Schapire (1997); Maillard & Munos (2010); Krichene et al. (2015) which enjoy $\mathcal{O}(d^{3/2}\sqrt{T})$ and $\tilde{\mathcal{O}}(\sqrt{dT})$ regret, respectively. We explicitly give several example algorithms and their regret guarantees in Appendix D. The primary drawback of these algorithms is that they are often computationally inefficient. In Section 6, we provide a more computationally efficient version of OSOCO at the cost of a slightly worse regret bound $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
4 Analysis
In this section, we provide regret and safety guarantees for OSOCO. In particular, Theorem 4.1 shows that, with high probability, OSOCO ensures the constraints are always satisfied and that the regret is $\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$ where $C_{\mathcal{A}}$ is the regret constant of the OCO algorithm $\mathcal{A}$ that is used by OSOCO (see Definition 3.2). It immediately follows that the regret of OSOCO is $\tilde{\mathcal{O}}(d\sqrt{T})$ when $C_{\mathcal{A}}$ is $\tilde{\mathcal{O}}(\sqrt{d})$, which holds when algorithms such as Hedge Freund & Schapire (1997) are used for $\mathcal{A}$ (as given in Appendix D).
Theorem 4.1.
Let $C_{\mathcal{A}}$ be the regret constant of $\mathcal{A}$ as defined in Definition 3.2. Fix some $\delta\in(0,1/2)$ and choose $\lambda=\max(1,D^{2})$, and $\beta_{t}$ according to Theorem 3.1. Then if $T\geq 3$, with probability at least $1-2\delta$, the regret of OSOCO (Algorithm 1) satisfies
$R_{T}\leq 4\frac{DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
and it holds that $Ax_{t}\leq b$ for all $t\in[T]$.
The proof of Theorem 4.1 (given in Appendix A) relies on a decomposition of the regret as
$R_{T}$ $=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\underbrace{\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})}_{\text{Cost of Safety}}+\underbrace{\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})}_{\text{Optimistic Regret}}$
We describe the analysis for the cost of safety in Section 4.1 and the optimistic regret in Section 4.2.
4.1 Cost of Safety
In this section, we study the cost that OSOCO incurs by ensuring constraint satisfaction, which we call the cost of safety. Specifically, the following lemma bounds the cumulative difference in cost between the played actions $x_{t}$ and the optimistic actions $\tilde{x}_{t}$. This is precisely the cost incurred by the algorithm scaling the optimistic action into the pessimistic set (as done in lines 10-11 in Algorithm 1).
Lemma 4.2.
Consider the same setup as Theorem 4.1. Then, it holds that
$\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})\leq 4\frac{DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}.$ (2)
Proof Sketch
The first step in the proof of Lemma 4.2 is to lower bound $\gamma_{t}$. Specifically, given any point $x\in\mathcal{Y}_{j}^{o}$, we find a scaling $\alpha\in[0,1]$ such that the scaled point $\alpha x$ is in $\mathcal{Y}_{j}^{p}$. Using this approach, it can be shown (Lemma A.2) that
$\gamma_{t}\geq 1-\frac{2}{b_{\mathrm{min}}}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$
where $j_{t}$ is the phase that round $t$ belongs to. Then, using the fact that all $f_{t}$ have bounded gradients, the lefthand side of (2) can be bounded with a constant multiple of $\sum_{t=1}^{T}(1-\gamma_{t})$ and therefore with a constant multiple of $\sum_{t=1}^{T}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}}$. From the phase termination criteria (line 17), the determinant of $V_{t}$ is no more than twice as large as the determinant of $\bar{V}_{j_{t}}$, so $\|x_{t}\|_{\bar{V}_{j_{t}}}^{-1}\leq 2\|x_{t}\|_{V_{t}}^{-1}$ (where we implicitly used Lemma 12 in Abbasi-Yadkori et al. (2011)). Also, $\beta_{t}$ is monotone in $t$ so it holds that $\beta_{T}\geq\bar{\beta}_{j_{t}}$ Putting these together, we get a bound in terms of $\beta_{T}\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}$ and therefore we can apply Cauchy-Schwarz and the elliptic potential lemma Dani et al. (2008); Abbasi-Yadkori et al. (2011) to get the result.
4.2 Optimistic Regret
In this section, we study the regret incurred by the optimistic actions, which we call the optimistic regret. This measure of regret is “optimistic” in that it is the regret that the algorithm would incur by using an optimistic estimate of the constraint.
Lemma 4.3.
Consider the same setup as Theorem 4.1. Then, it holds that
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})\leq C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
with probability at least $1-2\delta$.
Proof Sketch
Let $N$ be the number of phases, $t_{j}$ be the first round in the $j$th phase (where $t_{N+1}:=T+1$), and $x^{\star}_{j}$ be an optimal action in phase $j$, i.e.
$x^{\star}_{j}\in\operatorname*{arg\,min}_{x\in\mathcal{Y}_{j}^{o}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)$
It follows from Theorem 3.1 that $\mathcal{Y}$ is a subset of $\mathcal{Y}_{j}^{o}$ for all $j$ with probability at least $1-\delta$. As such, it holds with the same probability that $\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})\leq\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star})$, and therefore that
$\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star})$
$\leq\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})$
$\leq\sum_{j=1}^{N}C_{\mathcal{A}}\sqrt{t_{j+1}-t_{j}}$
$\leq C_{\mathcal{A}}\sqrt{NT},$
where the second inequality is the regret bound of $\mathcal{A}$ and the third inequality is Cauchy-Schwarz. Then, by analyzing the growth of the determinant of $V_{t}$, it can be shown that $N\leq 2d\log(T)$ (Lemma A.7). Finally, applying the Azuma inequality and union bound gives the result.
5 Zero Expected Constraint Violation
Thus far, we have viewed the constraint as being a fixed linear function ($Ax$) that the player gets noisy feedback on ($y_{t}$) and must satisfy at all rounds (at least with high probability). In this section, we take the alternative viewpoint that the constraint feedback $y_{t}$ is a stochastic function that must always be satisfied in expectation. We study this setting in Proposition 5.1 (proven in Appendix B), where we modify the OSOCO algorithm by slightly tightening the constraint, which ensures that $\mathbb{E}[y_{t}]\leq 0$ for all $t$ while only adding constants to the regret bound.
This slight modification of the problem setting allows for comparison with the literature on OCO under time-varying constraints, e.g. Neely & Yu (2017); Cao et al. (2018); Yi et al. (2021); Guo et al. (2022), and, in particular, OCO with stochastic constraints Yu et al. (2017); Castiglioni et al. (2022). Specifically, algorithms were given by Yu et al. (2017), Wei et al. (2020) and Castiglioni et al. (2022) showing that it is possible to guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ expected regret and $\tilde{\mathcal{O}}(\sqrt{T})$ expected violation when the constraints are iid and bounded convex functions, and the player receives full-information feedback on the constraints. We contribute to this literature by showing that it is possible to guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ expected regret and zero expected constraint violation when the constraint functions are linear in expectation and conditionally subgaussian, and the player receives bandit feedback on the constraints.
Proposition 5.1.
Consider the same setup as Theorem 4.1 except with OSOCO modified such that $b\leftarrow b-\mu\mathbf{1}$ for some $\mu\in[0,b_{\mathrm{min}})$. Then, it holds that $\mathbb{E}[y_{t}]\leq b$ for all $t\in[T]$ and
$\mathbb{E}[R_{T}]\leq\frac{8DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}+GD+\frac{b_{\mathrm{min}}G}{S}$
when $\mu=b_{\mathrm{min}}/T$ and $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$.
6 An Efficient Version
One drawback of the original OSOCO algorithm is that it requires the use of an online nonconvex optimization algorithm, which often do not allow for computationally efficient implementations. In this section, we modify the OSOCO algorithm to improve its computational efficiency, while only slightly weakening the regret guarantees to $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$. The key idea is that we modify the optimistic action set so that it can be represented as the union of a finite number of convex sets. We then give an efficient online optimization algorithm (to be used for $\mathcal{A}$) that is designed for this type of structured nonconvex action set.
6.1 Relaxation of Optimistic Action Set
In order for the optimistic action set to be amenable to efficient online optimization, we use the idea from Dani et al. (2008) of relaxing the least-squares confidence set. In particular, from the equivalence of norms, the optimistic action set can be relaxed to
$\mathcal{Y}_{j}^{o}=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\sqrt{d}\bar{\beta}_{j}\|\bar{V}_{j}^{-1/2}x\|_{\infty}\leq b\right\}.$
Although this set is still nonconvex, it can now be represented as the union of a finite number of convex sets. To explicitly define these convex sets, we use the idea from Cassel et al. (2022) of parameterizing the infinity norm by the active element in the vector and the sign of this element, i.e. $\|z\|_{\infty}=\max_{k\in[d],\xi\in\{-1,1\}}\xi z_{k}$ for a vector $z$. As such we can parameterize the optimistic set as
$\tilde{\mathcal{Y}}_{j}^{o}(k,\xi):=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\sqrt{d}\bar{\beta}_{j}\xi\mathbf{1}(\bar{V}_{j}^{-1/2})_{k}x\leq b\right\}.$
Evidently, $\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)$ is convex for any fixed $k,\xi\in[d]\times\{-1,1\}$ and furthermore the union over $k$ and $\xi$ is the original set, i.e. $\bigcup_{k,\xi\in[d]\times\{-1,1\}}\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)=\mathcal{Y}_{j}^{o}$. Accordingly, we modify OSOCO such that it sends the family of sets $\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$ to $\mathcal{A}$ in line 7. The regret guarantees for this modified algorithm are given Proposition 6.1 which are $\tilde{\mathcal{O}}((d^{3/2}+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$. The proof of Proposition 6.1 is given in Appendix C.
Proposition 6.1.
Consider the same setup as Theorem 4.1 except with OSOCO modified such that $\mathcal{Y}_{j}^{o}\leftarrow\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$. Then with probability at least $1-2\delta$,
$R_{T}\leq 8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}d\sqrt{2T\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
and $Ax_{t}\leq b$ for all $t\in[T]$.
6.2 HedgeDescent Algorithm
Algorithm 2 HedgeDescent
0:  $\{\mathcal{X}_{m}\}_{m\in[M]}$.
1:  Initialize: $x_{t}(m)=\mathbf{0}$, $p_{t}(m)=1/M$.
2:  for $t=1$ to $T$ do
3:     Sample Hedge: $m_{t}\sim p_{t}$.
4:     Play $x_{t}(m_{t})$ and observe $f_{t}$.
5:     Update experts: $x_{t+1}(m)=\Pi_{\mathcal{X}_{m}}\left(x_{t}(m)-\eta_{t}\nabla f_{t}(x_{t}(m))\right)$.
6:     Update Hedge: $p_{t+1}(m)\propto p_{t}(m)\exp(-\zeta_{t}f_{t}(x_{t}(m)))$.
7:  end for
In order to appropriately exploit the structure of the optimistic action set, we propose the online algorithm HedgeDescent (Algorithm 2) to be used for $\mathcal{A}$ in OSOCO. Since the optimistic action set is now the union of a finite number of convex sets, HedgeDescent uses a finite online algorithm to choose which convex set to play in and a continuous online algorithm to ensure low regret within each convex set. [^4] This approach is inspired by Cassel et al. (2022) where a similar idea is used to handle nonconvexity in the cost rather than nonconvexity in the constraint. More specifically, Hedge Freund & Schapire (1997) is used for the finite algorithm and online gradient descent Zinkevich (2003) is used for the continuous algorithm. As such, HedgeDescent only requires $M$ convex projections (where $M$ is the number of convex sets) and sampling from a distribution with support of cardinality $M$ in each round. When used with the modified OSOCO algorithm described in Proposition 6.1, $M=2d$ and therefore the combined algorithm requires $2d$ convex projections in each round.
We give the regret guarantee of HedgeDescent in Proposition 6.2 (proven in Appendix C) which is $\mathcal{O}(\sqrt{T\log(M)})$. Together with Proposition 6.1, it follows that the modified OSOCO algorithm enjoys $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret, which is only a factor of $\sqrt{d}$ worse than the regret guarantees for the original OSOCO algorithm.
Proposition 6.2.
Consider the online optimization setting described in Definition 3.2 except with action set $\bar{\mathcal{X}}=\bigcup_{m\in[M]}\mathcal{X}_{m}$ where each $\mathcal{X}_{m}$ is convex. If HedgeDescent (Algorithm 2) is played in this setting with $\zeta_{t}=\frac{\sqrt{4\log(M)}}{GD\sqrt{t}}$ and $\eta_{t}=\frac{D}{G\sqrt{t}}$, then it holds that
$\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\min_{x\in\bar{\mathcal{X}}}\sum_{t=1}^{T}f_{t}(x)\leq DG\sqrt{T\log(M)}+3DG\sqrt{T}.$
7 Numerical Experiments
(a) Safe online linear programming.
(b) Safe online quadratic programming.
Figure 1: Simulation results for our algorithm OSOCO and SO-PGD from Chaudhary & Kalathil (2022). The points indicate the average over 30 trials and the error bars are the standard deviation.
In order to verify the theoretical results and assess the empirical performance of OSOCO, we perform numerical experiments in several toy settings. As shown in Figure 1, we study the performance of OSOCO and SO-PGD Chaudhary & Kalathil (2022) in a safe online linear programming setting and a safe online quadratic programming setting. In the following, we give further details on these experiments and describe some additional results given in Appendix E.
Safe Online Linear Programming
This is a special case of our original problem where the cost functions are linear, i.e. $f_{t}(x)=\theta_{t}^{\top}x$. In particular, we consider a $2$-dimensional setting with $\theta_{t}\sim[0,1]^{d}$, $\mathcal{X}=\mathbb{B}$, $A=[I\ -I]^{\top}$, $b=(0.6)\mathbf{1}$ and $\epsilon_{t}\sim\mathcal{N}(0,\sigma I)$ where $\sigma=0.01$. We simulate OSOCO and SO-PGD for 30 realizations of this setting with algorithm parameters specified in Appendix E. The results are given in Figure 0(a), where the points show the average over the trials and the error bars show the standard deviation. This plot shows that the regret of OSOCO goes to constant $\sqrt{T}$ while having a smaller regret than the regret of SO-PGD for all durations considered.
Safe Online Quadratic Programming
This is the special case where the cost functions are quadratic. In particular, we consider quadratic functions of the form $f_{t}(x)=2\|x-v_{t}\|^{2}$. We consider a $2$-dimensional setting with $v_{t}\sim[-1,0]^{d}$, $\mathcal{X}=\mathbb{B}$, $A=[I\ -I]^{\top}$, $b=(0.5)\mathbf{1}$, and $\epsilon_{t}\sim\mathcal{N}(0,\sigma I)$ where $\sigma=0.01$. We simulate OSOCO and SO-PGD for 30 realizations of this setting with algorithm parameters given in Appendix E. The results are shown in Figure 0(b), with the points indicating the average of the trials and the error bars indicating the standard deviation. These results show that the regret of OSOCO goes to constant $\sqrt{T}$ and stays below SO-PGD for all horizon lengths considered.
Additional Results
In Appendix E, we give additional results comparing OSOCO to OGD Zinkevich (2003) and to the algorithm in Yu et al. (2017) which are given in Figure 2 and Figure 3, respectively. In Figure 4, we give some empirical evidence that the modified version of OSOCO given in Proposition 5.1 satisfies the constraint in expectation.
8 Conclusion
In this work, we give an algorithm with $\tilde{\mathcal{O}}(\sqrt{T})$ regret for the problem of OCO under unknown linear constraints and noisy constraint feedback. We then extend this algorithm to the setting of OCO with stochastic linear constraints and give a computationally efficient version. Numerical experiments verify the theoretical claims and provide an empirical comparison with baseline methods. One direction for future work is to investigate if our approach can be extended to the case where the adversary is non-oblivious and therefore can use past noise terms and actions when choosing cost functions.
Appendix A Proof of Theorem 4.1 (Regret of OSOCO)
In this section, we give the proof of Theorem 4.1. We first give the notation and assumptions in Section A.1 and state the high probability events in Section A.2. Then, we give the proof of Lemma 4.2 (Cost of Safety) in Section A.3 and the proof of Lemma 4.3 (Optimistic Regret) in Section A.4. We then complete the proof of Theorem 4.1 (Regret of OSOCO) in Section A.5. Lastly, side lemmas are given in Section A.6.
A.1 Notation and Assumptions
In addition to the notation given in the body of the paper, we use $N$ to refer to the total number of phases and $t_{j}$ to refer to the first round within phase $j$ where $t_{N+1}:=T$. Additionally, we use $j_{t}$ to refer to the phase to which round $t$ belongs, i.e. $j_{t}=\max\{j\in[N]:t_{j}\leq t\}$.
In addition to the assumptions made in the body, we will also assume throughout this section that the assumptions made in Theorem 4.1 hold, i.e. $\lambda=\max(1,D^{2})$, $\beta_{t}:=\rho\sqrt{d\log\left(\frac{1+(t-1)L^{2}/\lambda}{\delta/n}\right)}$, that $\mathcal{A}$ enjoys a regret constant $C_{\mathcal{A}}$ (in the sense of Definition 3.2) and that $T\geq 3$.
A.2 High probability events
In this section, we show that with an appropriate choice of problem parameters, the events
$\mathcal{E}_{\mathrm{conf}}:=\left\{(A-\hat{A}_{j})x\in\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbb{B}_{\infty},\forall x\in\mathbb{R}^{d},\forall j\in[N]\right\},$
and
$\mathcal{E}_{\mathrm{azuma}}:=\left\{\sum_{t=1}^{T}(f_{t}(\tilde{x}_{t})-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})])\leq 2DG\sqrt{2T\log(1/\delta)}\right\},$
occur with high probability. First, it follows from Theorem 3.1 that $\mathbb{P}(\mathcal{E}_{\mathrm{conf}})\geq 1-\delta$. Then, an application of Azuma’s inequality yields a bound on the probability of $\mathcal{E}_{\mathrm{azuma}}$ as shown in the following lemma.
Lemma A.1 (Probability of $\mathcal{E}_{\mathrm{azuma}}$).
The event $\mathcal{E}_{\mathrm{azuma}}$ occurs with probability at least $1-\delta$.
Proof.
Define $Y_{t}=f_{t}(\tilde{x}_{t})-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]$ and note that $Y_{t}$ is a martingale difference sequence since $\mathbb{E}_{t}[Y_{t}]=0$ and $\mathbb{E}_{t}[Y_{t-1}]=Y_{t-1}$. Furthermore, we can fix some $y\in\mathcal{X}$ and then it follows from Assumption 2.1 and 2.2 that
$\begin{split}|Y_{t}|&=|f_{t}(\tilde{x}_{t})-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]|\\ &=|f_{t}(\tilde{x}_{t})-f_{t}(y)+f_{t}(y)-\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]|\\ &=|f_{t}(\tilde{x}_{t})-f_{t}(y)+\mathbb{E}_{t}[f_{t}(y)-f_{t}(\tilde{x}_{t})]|\\ &\leq|f_{t}(\tilde{x}_{t})-f_{t}(y)|+\mathbb{E}_{t}[|f_{t}(y)-f_{t}(\tilde{x}_{t})|]\\ &\leq 2G\|\tilde{x}_{t}-y\|\\ &\leq 2GD.\end{split}$ (3)
Then, applying Azuma’s inequality ensures that
$\sum_{t=1}^{T}Y_{t}\leq 2DG\sqrt{2T\log(1/\delta)}$
with probability at least $1-\delta$. ∎
A.3 Proof of Lemma 4.2 (Cost of Safety)
Before proving Lemma 4.2, we first give a key lemma lower bounding $\gamma_{t}$.
Lemma A.2 (Bound on $\gamma_{t}$).
It holds for all $t\in[T]$ that
$\gamma_{t}\geq 1-\frac{2}{b_{\mathrm{min}}}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}.$
Proof.
First, we study the scaling that is required to take any $x\in\mathcal{Y}_{j}^{o}$ into $\mathcal{Y}_{j}^{p}$. It holds (by definition) that
$\hat{a}_{j,i}^{\top}x-\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\leq b_{i}\quad\Longleftrightarrow\quad\hat{a}_{j,i}^{\top}x\leq b_{i}+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}.$
Then, choosing $\alpha_{i}=\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}}$ such that $b_{i}\alpha_{i}=b_{i}-2\alpha_{i}\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}$ and $\alpha_{i}\in[0,1]$, we get that
$\hat{a}_{j,i}^{\top}(\alpha_{i}x)\leq\alpha_{i}\left(b_{i}+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\right)=\alpha_{i}b_{i}+\alpha_{i}\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}=b_{i}-\bar{\beta}_{j}\|(\alpha_{i}x)\|_{\bar{V}_{j}^{-1}}.$
This means that $\alpha_{i}x$ satisfies the $i$th constraint of the pessimistic set. Therefore, we can ensure that it holds for all constraints by choosing
$\bar{\alpha}:=\min_{i\in[n]}\left(\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}}\right)=\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}},$
where we use the fact that $\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}}$ is monotone in $b_{i}$. Since $\bar{\alpha}\in[0,1]$, $\mathbf{0}\in\mathcal{X}$, and $\mathcal{X}$ is convex, it holds that $\bar{\alpha}x\in\mathcal{Y}^{p}_{j}$. Because $\tilde{x}_{t}$ is in $\mathcal{Y}_{j_{t}}^{o}$, we can apply $x\leftrightarrow\tilde{x}_{t}$ to get that $\bar{\alpha}\tilde{x}_{t}$ is in $\mathcal{Y}^{p}_{j_{t}}$. Since $\gamma_{t}$ is the largest such scaling, $\bar{\alpha}$ is a lower bound on $\gamma_{t}$, i.e.
$\gamma_{t}=\max\{\alpha\in[0,1]:\alpha\tilde{x}_{t}\in\mathcal{Y}_{j_{t}}^{p}\}\geq\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j_{t}}\|\tilde{x}_{t}\|_{\bar{V}_{j_{t}}^{-1}}}.$ (4)
Since, $x_{t}=\gamma_{t}\tilde{x}_{t}$ and $\gamma_{t}\geq 0$, it follows that $\gamma_{t}\|\tilde{x}_{t}\|_{\bar{V}_{j_{t}}^{-1}}=\|\gamma_{t}\tilde{x}_{t}\|_{\bar{V}_{j_{t}}^{-1}}=\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$. Therefore, we can rearrange (4) to get that
$\gamma_{t}\geq 1-2\frac{1}{b_{\mathrm{min}}}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}},$
completing the proof. ∎
With this, we then restate and prove Lemma 4.2.
Lemma A.3 (Duplicate of Lemma 4.2).
It holds that
$\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}.$ (5)
Proof.
We can study the lefthand side of (5) to get that
$\sum_{t=1}^{T}\left(f_{t}(\tilde{x}_{t})-f_{t}(x_{t})\right)$ $\leq G\sum_{t=1}^{T}\|\tilde{x}_{t}-x_{t}\|$
$\leq G\sum_{t=1}^{T}\|\tilde{x}_{t}-\gamma_{t}\tilde{x}_{t}\|$ (a)
$\leq DG\sum_{t=1}^{T}|1-\gamma_{t}|$
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\bar{\beta}_{j}\|x_{t}\|_{\bar{V}_{j}^{-1}}$ (b)
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\|x_{t}\|_{\bar{V}_{j}^{-1}}$ (c)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}$ (d)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{T\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}^{2}}$ (e)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(1+\frac{T}{\lambda d}\right)}$ (f)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)},$ (g)
where (a) uses the fact that $x_{t}=\gamma_{t}\tilde{x}_{t}$, (b) follows from Lemma A.2, (c) uses the fact that $\beta_{T}\geq\beta_{t}$ for all $t\in[T]$, (d) uses the fact that for pd matrices $B\preceq C$ it holds that $\|z\|_{B^{-1}}\leq\frac{\det(C)}{\det(B)}\|z\|_{C^{-1}}$ for any $z$ (see e.g. Lemma 12 in Abbasi-Yadkori et al. (2011) or Lemma 27 in Cohen et al. (2019)), (e) is Cauchy-Schwarz, (f) follows from the elliptic potential lemma (Lemma A.6) and (g) is from the fact that $2d\log\left(1+\frac{T}{\lambda d}\right)\leq 2d\log\left(1+T\right)\leq 3d\log\left(T\right)$ when $T\geq 3$ and $\lambda=\max(1,D^{2})$. ∎
A.4 Proof of Lemma 4.3 (Optimistic Regret)
We restate and prove Lemma 4.3 in the following.
Lemma A.4 (Duplicate of Lemma 4.3).
Conditioned on $\mathcal{E}_{\mathrm{conf}}$ and $\mathcal{E}_{\mathrm{azuma}}$, it holds that
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})\leq C_{\mathcal{A}}\sqrt{4dT\log(T)}+DG\sqrt{2T\log(1/\delta)}$
Proof.
Let $x^{\star}_{j}$ be an optimal action within the optimistic set in phase $j$, i.e. $x^{\star}_{j}\in\operatorname*{arg\,min}_{x\in\mathcal{Y}_{j}^{o}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)$. Then, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that $x^{\star}\in\mathcal{Y}\subseteq\mathcal{Y}_{j}^{o}$ and therefore the optimal cost in each phase lower bounds the cost of the optimal action in each phase, i.e.
$\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})\leq\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}).$ (6)
Then, to apply the regret guarantee of $\mathcal{A}$, we need that the expectation over the distribution chosen by $\mathcal{A}$ is equal to the expectation conditioned on the randomness in previous rounds. Indeed, the distribution chosen by $\mathcal{A}$ in round $t$ (denoted $p_{t}$) is fully determined by the randomness in previous rounds (and $f_{t}$ is fixed) so it holds that $\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]=\mathbb{E}_{\tilde{x}_{t}\sim p_{t}}[f_{t}(\tilde{x}_{t})]$, and therefore
$\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})=\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{\tilde{x}_{t}\sim p_{t}}[f_{t}(\tilde{x}_{t})]-\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})\leq C_{\mathcal{A}}\sqrt{t_{j+1}-t_{j}}$ (7)
for all $j\in[N]$. Also, note that the duration of each phase ($t_{j+1}-t_{j}$) depends on the actions chosen in that phase. Nonetheless, $\mathcal{A}$ is specified to handle a setting that can be terminated at any time and therefore the regret guarantees hold even when the duration of the phase depends on the actions taken.
Then, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that
$\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star})$
$\leq\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})$ (a)
$\leq\sum_{j=1}^{N}C_{\mathcal{A}}\sqrt{t_{j+1}-t_{j}}$ (b)
$\leq C_{\mathcal{A}}\sum_{j=1}^{N}\sqrt{t_{j+1}-t_{j}}$ (c)
$\leq C_{\mathcal{A}}\sqrt{N\sum_{j=1}^{N}(t_{j+1}-t_{j})}$ (d)
$=C_{\mathcal{A}}\sqrt{NT}$
$\leq C_{\mathcal{A}}\sqrt{4d\log(T)T}$ (e)
where (a) uses (6), (b) uses (7), (c) is from the specification that $\mathcal{A}$ is polylogarithmic in $t_{j+1}-t_{j}$ and therefore (with abuse of notation) pull it out of the summation to mean that it is polylogarithmic in $T$ as $t_{j+1}-t_{j}\leq T$, (d) is Cauchy-Schwarz and (e) uses $N\leq 4d\log(T)$ from Lemma A.7.
Then, also conditioning on $\mathcal{E}_{\mathrm{azuma}}$, it holds
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]+\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=2DG\sqrt{2T\log(1/\delta)}+C_{\mathcal{A}}\sqrt{4d\log(T)T},$
completing the proof. ∎
A.5 Completing the proof
We restate and prove Theorem 4.1 in the following.
Theorem A.5 (Duplicate of Theorem 4.1).
Fix some $\delta\in(0,1/2)$. Then, with probability at least $1-2\delta$, the regret of OSOCO (Algorithm 1) satisfies
$R_{T}\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
and $Ax_{t}\leq b$ for all $t\in[T]$.
Proof.
We can apply Lemmas 4.2 and 4.3 and condition on $\mathcal{E}_{\mathrm{conf}}$ and $\mathcal{E}_{\mathrm{azuma}}$ to get that
$R_{T}$ $=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})+\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+2DG\sqrt{2T\log(1/\delta)}+C_{\mathcal{A}}\sqrt{4d\log(T)T}.$
Also, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that $x_{t}\in\mathcal{Y}_{j_{t}}^{p}\subseteq\mathcal{Y}$ for all $t\in[T]$ and therefore $Ax_{t}\leq b$ for all $t\in[T]$. To get a bound of the probability of both events holding, we use Theorem 3.1 and Lemma A.1 with the union bound to get that
$\mathbb{P}(\mathcal{E}_{\mathrm{conf}}\cap\mathcal{E}_{\mathrm{azuma}})=1-\mathbb{P}(\mathcal{E}_{\mathrm{conf}}^{c}\cup\mathcal{E}_{\mathrm{azuma}}^{c})\geq 1-(\mathbb{P}(\mathcal{E}_{\mathrm{conf}}^{c})+\mathbb{P}(\mathcal{E}_{\mathrm{azuma}}^{c}))\geq 1-2\delta,$
completing the proof. ∎
A.6 Side lemmas
In this section, we give side lemmas that were used in the proof. First, we have the well-known elliptic potential lemma. The version that we give is from Abbasi-Yadkori et al. (2011).
Lemma A.6 (Elliptic Potential, Lemma 11 in Abbasi-Yadkori et al. (2011)).
Consider a sequence $(w_{t})_{t\in\mathbb{N}}$ where $w_{t}\in\mathbb{R}^{d}$ and $\|w_{t}\|\leq L$ for all $t\in\mathbb{N}$. Let $W_{t}=\lambda I+\sum_{s=1}^{t-1}w_{s}w_{s}^{\top}$ for some $\lambda\geq\max(1,L^{2})$. Then, it holds for any $m\in\mathbb{N}$ that
$\sum_{t=1}^{m}\|w_{t}\|_{W_{t}^{-1}}^{2}\leq 2d\log\left(1+\frac{m}{\lambda d}\right)$
Then, we give a bound on the number of phases in the following lemma.
Lemma A.7 (Number of Phases).
It holds that $N\leq 4d\log(T)$.
Proof.
Because $\det(\bar{V}_{j+1})\geq 2\det(\bar{V}_{j})$ for all $j\in[1,N-1]$ by definition, it holds that $\det(V_{T})\geq\det(\bar{V}_{N})\geq 2^{N-1}\det(\bar{V}_{1})$. Then, taking the log of both sides and using the fact that $N\leq 1+2\log(2^{N-1})$, we have
$N$ $\leq 1+2\log\left(\frac{\det(V_{T})}{\det(\bar{V}_{1})}\right)$
$=1+2\log\det(\bar{V}_{1}^{-1/2}V_{T}\bar{V}_{1}^{-1/2})$
$\leq 1+2d\log\|\bar{V}_{1}^{-1/2}V_{T}\bar{V}_{1}^{-1/2}\|$
$=1+2d\log\left\|\bar{V}_{1}^{-1/2}\left(\sum_{t=1}^{T}x_{t}x_{t}^{\top}+\bar{V}_{1}\right)\bar{V}_{1}^{-1/2}\right\|$
$=1+2d\log\left\|\frac{1}{\lambda}\sum_{t=1}^{T}x_{t}x_{t}^{\top}+I\right\|$
$\leq 1+2d\log\left(\frac{1}{\lambda}\sum_{t=1}^{T}\left\|x_{t}x_{t}^{\top}\right\|+1\right)$
$=1+2d\log\left(\frac{1}{\lambda}\sum_{t=1}^{T}\left\|x_{t}\right\|^{2}+1\right)$
$\leq 1+2d\log\left(\frac{TD^{2}}{\lambda}+1\right)$
$\leq 1+2d\log\left(T+1\right)$
$\leq 4d\log\left(T\right),$
where the second-to-last inequality uses $\lambda\geq D^{2}$ and the last inequality uses $T\geq 3$. ∎
Appendix B Proof of Proposition 5.1 (Zero Expected Violation)
In this section, we restate and prove Proposition 5.1.
Proposition B.1 (Restatement of Proposition 5.1).
Consider the same setup as Theorem 4.1 except with $b\leftarrow b-\mu\mathbf{1}$ for $\mu\in[0,b_{\mathrm{min}})$. Then, it holds that $\mathbb{E}[y_{t}]\leq b$ for all $t\in[T]$ and
$\mathbb{E}[R_{T}]\leq 8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(SDT/b_{\mathrm{min}})}+GD+\frac{b_{\mathrm{min}}G}{S}$
when $\mu=b_{\mathrm{min}}/T$ and $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$.
Proof.
Let the event that Theorem 4.1 holds be denoted by $\mathcal{E}$ which occurs with probability at least $1-2\delta$. Then, running OSOCO with $b\leftrightarrow b-\mu\mathbf{1}$ ensures that $Ax_{t}\leq b-\mu\mathbf{1}$ for all $t\in[T]$ conditioned on $\mathcal{E}$. Therefore,
$\mathbb{E}[y_{t,i}]$ $=\mathbb{E}[a_{i}^{\top}x_{t}+\epsilon_{t,i}]$
$=\mathbb{E}[a_{i}^{\top}x_{t}]$
$=\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}\}]+\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}^{c}\}]$
$\leq b_{i}-\mu+2SD\delta$
$\leq b_{i}-b_{\mathrm{min}}/T+b_{\mathrm{min}}/T$
$=b_{i}$
where we use the trivial constraint bound $a_{i}^{\top}x_{t}\leq SD$ in the first inequality and then the choice of $\mu$ and $\delta$ in the last step. This gives the safety guarantee.
Next, we study the impact on the regret. Let $\tilde{\mathcal{Y}}=\{x\in\mathcal{X}:Ax\leq b-\mathbf{1}\mu\}$. First, note that when $\mu\in[0,b_{\mathrm{min}}]$, it holds that $\alpha x^{\star}$ is in $\tilde{\mathcal{Y}}$ where $\alpha=\frac{b_{\mathrm{min}}-\mu}{b_{\mathrm{min}}}$ as
$a_{i}^{\top}(\alpha x^{\star})=\alpha a_{i}^{\top}x^{\star}\leq\alpha b_{i}=b_{i}\frac{b_{\mathrm{min}}-\mu}{b_{\mathrm{min}}}\leq b_{i}\frac{b_{i}-\mu}{b_{i}}=b_{i}-\mu$
for all $i\in[n]$, where we use the fact that the form $\frac{y-\mu}{y}$ is increasing in $y$ when $y>0$. Therefore, we know that
$\|\Pi_{\tilde{\mathcal{Y}}}(x^{\star})-x^{\star}\|\leq\|\alpha x^{\star}-x^{\star}\|=(1-\alpha)\|x^{\star}\|\leq\frac{\mu}{b_{\mathrm{min}}}D.$
Finally, with $\bar{x}^{\star}=\operatorname*{arg\,min}_{x\in\tilde{\mathcal{Y}}}\sum_{t=1}^{T}f_{t}(x)$ and $R_{T}^{\mathrm{osoco}}$ as the (high probability) regret bound of OSOCO, it holds conditioned on $\mathcal{E}$ that
$R_{T}=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\bar{x}^{\star})+\sum_{t=1}^{T}f_{t}(\bar{x}^{\star})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq R_{T}^{\mathrm{osoco}}+\sum_{t=1}^{T}f_{t}(\bar{x}^{\star})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq R_{T}^{\mathrm{osoco}}+\sum_{t=1}^{T}f_{t}(\Pi_{\tilde{\mathcal{Y}}}(x^{\star}))-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq R_{T}^{\mathrm{osoco}}+G\sum_{t=1}^{T}\|\Pi_{\tilde{\mathcal{Y}}}(x^{\star})-x^{\star}\|$
$\leq R_{T}^{\mathrm{osoco}}+GDT\frac{\mu}{b_{\mathrm{min}}}$
$\leq R_{T}^{\mathrm{osoco}}+GD$
where we use the choice of $\mu$ in the last step. However, we care about the expected regret, so we get
$\mathbb{E}[R_{T}]$ $=\mathbb{E}[R_{T}\mathbb{I}\{\mathcal{E}\}]+\mathbb{E}[R_{T}\mathbb{I}\{\mathcal{E}^{c}\}]$
$\leq R_{T}^{\mathrm{osoco}}+GD+2GDT\delta$
$\leq R_{T}^{\mathrm{osoco}}+GD+\frac{b_{\mathrm{min}}G}{S},$
where we use the trivial bound on the regret given by
$R_{T}=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})\leq G\sum_{t=1}^{T}\|x_{t}-x^{\star}\|\leq GDT.$
Appendix C Proof of Propositions 6.1 and 6.2 (Improved Efficiency)
We give proofs of Propositions 6.1 (Relaxed OSOCO) and 6.2 (HedgeDescent) in Sections C.1 and C.2, respectively.
C.1 Proof of Proposition 6.1 (Relaxed OSOCO)
We restate and prove Proposition 6.1 in the following.
Proposition C.1 (Duplicate of Proposition 6.1).
Consider the same setup as Theorem 4.1 except with OSOCO modified such that $\mathcal{Y}_{j}^{o}\leftrightarrow\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$. Then, it holds with probability at least $1-2\delta$ that
$R_{T}\leq\mbox{ }$ $8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}d\sqrt{2T\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
and $Ax_{t}\leq b$ for all $t\in[T]$.
Proof.
First note that, following the same process as Lemma A.2, it holds that $\gamma_{t}\geq 1-\frac{1}{b_{\mathrm{min}}}\sqrt{d}\bar{\beta}_{j_{t}}\|\bar{V}_{j_{t}}^{-1/2}x_{t}\|_{\infty}$. Then, using the equivalence of norms, it follows that
$\gamma_{t}\geq 1-\frac{1}{b_{\mathrm{min}}}\sqrt{d}\bar{\beta}_{j_{t}}\|\bar{V}_{j_{t}}^{-1/2}x_{t}\|_{\infty}\geq 1-\frac{1}{b_{\mathrm{min}}}\sqrt{d}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$
Following the rest of the proof of Theorem 4.1 with this bound gives the result. ∎
C.2 Proof of Proposition 6.2 (HedgeDescent)
We restate and prove Proposition 6.2 in the following.
Proposition C.2 (Duplicate of Proposition 6.2).
Consider an OCO setting with action set $\bar{\mathcal{X}}=\bigcup_{m\in[M]}\mathcal{X}_{m}$ where each $\mathcal{X}_{m}$ is convex and $\|\nabla f_{t}(x)\|\leq G$ and $\|x\|\leq D/2$ for all $x\in\bar{\mathcal{X}}$ and $t\in[T]$ (as in Definition 3.2). If HedgeDescent (Algorithm 2) is played in this setting with $\zeta_{t}=\frac{\sqrt{4\log(M)}}{GD\sqrt{t}}$ and $\eta_{t}=\frac{D}{G\sqrt{t}}$, then it holds that
$\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\min_{x\in\bar{\mathcal{X}}}\sum_{t=1}^{T}f_{t}(x)\leq DG\sqrt{T\log(M)}+3DG\sqrt{T}.$
Proof.
First, note that we can apply the standard regret guarantees of online gradient descent (e.g. Theorem 3.1 Hazan et al. (2016)) to get that the regret of each expert $m\in[M]$ is
$\sum_{t=1}^{T}f_{t}(x_{t}(m))-\sum_{t=1}^{T}f_{t}(x^{\star}_{m})\leq 3DG\sqrt{T},$
where $x^{\star}_{m}=\operatorname*{arg\,min}_{x\in\mathcal{X}_{m}}\sum_{t=1}^{T}f_{t}(x)$. Then, we cast the problem of choosing $m$ as an expert advice problem. To put this into the standard form, we shift and scale the losses by defining the surrogate losses $\ell_{t}(m)=\frac{f_{t}(x_{t}(m))-\bar{f}_{t}}{GD}$ where $\bar{f}_{t}=\min_{y\in\bar{\mathcal{X}}}f_{t}(y)$. This ensures that the $\ell_{t}(m)\in[0,1]$. Then, the Hedge update becomes
$p_{t+1}(m)$ $\propto p_{t}(m)\exp(-\zeta_{t}f_{t}(x_{t}(m)))$
$=p_{t}(m)\exp(-\zeta_{t}(GD\ell_{t}(m)+\bar{f}_{t}))$
$=p_{t}(m)\exp(-\zeta_{t}\bar{f}_{t})\exp(-\zeta_{t}GD\ell_{t}(m))$
$\propto p_{t}(m)\exp(-(\sqrt{4\log(M)}/\sqrt{t})\ell_{t}(m)).$
Note that this update is exactly the form of the classical Hedge algorithm with step size $\bar{\zeta}_{t}=\sqrt{4\log(M)}/\sqrt{t}$. As such, using the standard analysis of Hedge with time-dependent step size, e.g. Chernov & Zhdanov (2010), it holds that
$\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[\ell_{t}(m_{t})]-\sum_{t=1}^{T}\ell_{t}(m^{\star})\leq\sqrt{T\log(M)},$
where $m^{\star}$ is such that $x^{\star}\in\mathcal{X}_{m^{\star}}$. Putting everything together, we get that
$\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))+\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=DG\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[\ell_{t}(m_{t})]-DG\sum_{t=1}^{T}\ell_{t}(m^{\star})+\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq DG\sqrt{T\log(M)}+3DG\sqrt{T},$
where we use the fact that $\sum_{t=1}^{T}f_{t}(x^{\star})=\sum_{t=1}^{T}f_{t}(x^{\star}_{m^{\star}})$. ∎
Appendix D Example OCO Algorithms
In this section, we give examples of algorithms that can be used for the OCO algorithm $\mathcal{A}$ in OSOCO, i.e. those that admit the type of regret bound specified in Definition 3.2.
D.1 Problem setting
We consider a setting where at each round $t$, the player chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and, then the player observes the cost function $f_{t}$ and suffers the cost $f_{t}(x_{t})$. The game can be terminated at any time and therefore the final round $T$ is unknown to the player. In the algorithms below, we use the flag $\mathsf{alive}$ to indicate that the game has not been terminated. Also, $\mathcal{X}$ is not necessarily convex. We assume that all $f_{t}$ are differentiable and $\|\nabla f_{t}(x)\|\leq G$ for all $x\in\mathcal{X}$. We also assume that $\|x\|\leq D/2$ for all $x\in\mathcal{X}$. We also take the adversary to be oblivious such that the cost functions are fixed ahead of time.
D.2 Example Algorithm 1: Anytime Hedge over a Finite Cover
The simplest algorithm to handle this problem is Hedge played over a finite cover with the doubling trick. We give pseudocode for this algorithm in Algorithm 3. This is a well-known extension of the Hedge algorithm Freund & Schapire (1997) that has been noted by works such as Dani et al. (2007), Maillard & Munos (2010) and Krichene et al. (2015).
Algorithm 3 Anytime Hedge over a Finite Cover
0:  $\mathcal{X}$.
1:  Initialize: $t=1$, $j=0$, $\mathsf{alive}=\mathtt{True}$.
2:  while $\mathsf{alive}$ do
3:     Construct $\Delta_{j}$-net over $\mathcal{X}$, i.e. choose points $\{z_{1},...,z_{M_{j}}\}$ such that $\min_{m\in[M_{j}]}\|x-z_{m}\|\leq\Delta_{j}$ for all $x\in\mathcal{X}$.
4:     Set distribution $p_{t}$ over $[M_{j}]$ to be uniform.
5:     while $t\leq\tau_{j+1}-1$ and $\mathsf{alive}$ do
6:        Sample $m_{t}\sim p_{t}$.
7:        Play $x_{t}=z_{m_{t}}$ and observe $f_{t}$.
8:        Update $p_{t+1}(m)\propto\exp(-\eta_{j}f(z_{m}))$.
9:        $t=t+1$.
10:     end while
11:     $j=j+1$
12:  end while
Then, we give the regret guarantees of Algorithm 3.
Proposition D.1.
Consider Algorithm 3 with $\Delta_{j}=1/(G\sqrt{\tau_{j}})$ and $\eta_{j}=\sqrt{2\log(M_{j})/\tau_{j}}/GD$. Then, it holds that
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x)\leq 5\left(GD\sqrt{4d\log(2DG\sqrt{dT})}+2\right)\sqrt{T}$
Proof.
Let $x^{\star}$ be the optimal action over the entire time horizon, $x^{\star}_{j}$ be the optimal action within phase $j$ and $m^{\star}_{j}$ be the index of the nearest net point. Also, let $\tau_{j}$ be the duration of phase $j$ and $j=N$ be the index of the last phase. To put the update into the standard form for Hedge, we shift and scale the losses by defining the surrogate losses $\ell_{t}(m)=\frac{f_{t}(z_{m})-\bar{f}_{t}}{GD}$ where $\bar{f}_{t}=\min_{y\in\mathcal{X}}f_{t}(y)$. This ensures that the $\ell_{t}(m)\in[0,1]$. Then, the Hedge update becomes
$p_{t+1}(m)$ $\propto p_{t}(m)\exp(-\eta_{j}f_{t}(z_{m}))$
$=p_{t}(m)\exp(-\eta_{j}(GD\ell_{t}(m)+\bar{f}_{t}))$
$=p_{t}(m)\exp(-\eta_{j}\bar{f}_{t})\exp(-\eta_{j}GD\ell_{t}(m))$
$\propto p_{t}(m)\exp(-\eta_{j}GD\ell_{t}(m))$
$=p_{t}(m)\exp(-(\sqrt{2\log(M_{j})/\tau_{j}})\ell_{t}(m)).$
This is equivalent to the classical analysis of Hedge Arora et al. (2012) with step size $\eta_{j}^{\prime}=\sqrt{2\log(M_{j})/\tau_{j}}$ and therefore, for any phase $j\in[0,N]$,
$\begin{split}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{m\in M_{j}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m})&=GD\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[\ell_{t}(x_{t})]-GD\min_{m\in M_{j}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\ell_{t}(z_{m})\\ &\leq GD\sqrt{4\tau_{j}\log(M_{j})},\end{split}$ (8)
Note that if the game is terminated in the midst of the final round, there will not be cost functions for the remaining rounds. However, we can define the “fictitious” cost functions $f_{T+1},...,f_{\tau_{N+1}-1}$ that extend from the end of the game to the end of the final phase. This is why we have the regret guarantee for all phases as defined in (8).
Also, for any $x\in\mathcal{X}$, there exists $z\in\{z_{1},...,z_{M_{j}}\}$ such that
$f_{t}(x)-f_{t}(z)\leq G\|x-z\|\leq G\Delta_{j}$
Then, we get that within a given phase $j$, it holds that
$\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$ $\leq\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m^{\star}_{j}})+\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(z_{m^{\star}_{j}})-\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$
$\leq GD\sqrt{4\tau_{j}\log(M_{j})}+G\Delta_{j}\tau_{j}$
Furthermore, it is known (e.g. Example 27.1 in Shalev-Shwartz & Ben-David (2014)) that we can construct a grid for the net to get $M_{j}\leq(2D\sqrt{d}/\Delta_{j})^{d}$ for desired $\Delta_{j}$. Therefore, choosing $\Delta_{j}=1/(G\sqrt{\tau_{j}})$ will need $M_{j}\leq(2DG\sqrt{d\tau_{j}})^{d}$ points. Also, note that $N$ is the smallest integer such that $\sum_{j=0}^{N}2^{j}=2^{N+1}-1\geq T$ and therefore $N=\lceil\log_{2}(T+1)-1\rceil\leq\log_{2}(T+1)$. Therefore, it holds that
$\sum_{j=0}^{N}\sqrt{\tau_{j}}=\sum_{j=0}^{N}\sqrt{2^{j}}=\sum_{j=0}^{N}\sqrt{2}^{j}=\frac{1-\sqrt{2}^{N+1}}{1-\sqrt{2}}=\frac{\sqrt{2^{N+1}}-1}{\sqrt{2}-1}\leq 5\sqrt{T}$ (9)
Putting everything together, we get that
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star})$
$\leq\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$
$\leq\sum_{j=0}^{N}(GD\sqrt{4\tau_{j}\log(M_{j})}+2G\Delta_{j}\tau_{j})$
$=\sum_{j=0}^{N}\left(GD\sqrt{4d\tau_{j}\log(2DG\sqrt{d\tau_{j}})}+2\sqrt{\tau_{j}}\right)$
$\leq\left(GD\sqrt{4d\log(2DG\sqrt{dT})}+2\right)\sum_{j=0}^{N}\sqrt{\tau_{j}}$
$\leq 5\left(GD\sqrt{4d\log(2DG\sqrt{dT})}+2\right)\sqrt{T}$
completing the proof. ∎
D.3 Example Algorithm 2: Anytime Follow the Perturbed Leader
Another algorithm that can be used for this setting is Follow the Perturbed Leader (FTPL) Kalai & Vempala (2005) which has been popular for a variety of OCO settings. We use the version from Suggala & Netrapalli (2020) with an exact optimization oracle. We use the version and analysis from Suggala & Netrapalli (2020) because they study it in the nonconvex setting, which is directly applicable to our setting because the action set is nonconvex. We give the FTPL algorithm with the doubling trick in Algorithm 4.
Algorithm 4 Anytime Follow the Perturbed Leader
0:  $\mathcal{X}$.
1:  Initialize: $t=1$, $j=0$, $\mathsf{alive}=\mathtt{True}$, $\tau_{j}:=2^{j}$.
2:  while $\mathsf{alive}$ do
3:     while $t\leq\tau_{j+1}-1$ and $\mathsf{alive}$ do
4:        Let $\sigma_{t}=[\sigma_{t,1}\ ...\ \sigma_{t,d}]^{\top}$ where $\sigma_{t,i}\sim\mathrm{Exp}(\eta_{j})$.
5:        Choose $x_{t}\in\operatorname*{arg\,min}_{x\in\mathcal{X}}\left(\sum_{s=1}^{t-1}f_{t}(x)-\sigma_{t}^{\top}x\right)$
6:        Play $x_{t}$ and observe $f_{t}$.
7:        $t=t+1$.
8:     end while
9:     $j=j+1$
10:  end while
We then give the regret guarantees of this algorithm in the following.
Proposition D.2.
Consider Algorithm 4 with $\eta_{j}=1/\sqrt{2dG^{2}\tau_{j}}$. Then, it holds that
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x)\leq 630DGd^{3/2}\sqrt{2T},$
where $p_{t}$ is the distribution over the action round $t$.
Proof.
We will use the analysis from Suggala & Netrapalli (2020) to bound the regret within each phase. First note that Suggala & Netrapalli (2020) assumes that $\|x-y\|_{\infty}\leq D^{\prime}$ for all $x,y\in\mathcal{X}$ and that $f_{t}$ are $L^{\prime}$-Lipschitz w.r.t. the 1-norm. In our setting, it holds that for all $x,y\in\mathcal{X}$ that
$\|x-y\|_{\infty}\leq\|x-y\|=\|x-\mathbf{0}+\mathbf{0}-y\|\leq\|x-\mathbf{0}\|+\|\mathbf{0}-y\|\leq D$
and that
$|f_{t}(x)-f_{t}(y)|\leq G\|x-y\|\leq G\|x-y\|_{1}.$
Therefore, we can take $D^{\prime}=D$ and $L^{\prime}=G$. Also, Suggala & Netrapalli (2020) assume access to a $(\alpha,\beta)$ inexact oracle, but we assume that this oracle is exact and therefore take $\alpha,\beta=0$. Then, since the distribution over actions chosen by FTPL in each round (denoted by $p_{t}$) is independent of past randomness when the adversary is oblivious, we can apply Theorem 1 in Suggala & Netrapalli (2020) to get that
$\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x)$ $\leq 125\eta_{j}(L^{\prime})^{2}d^{2}D^{\prime}2\tau_{j}+\frac{dD^{\prime}}{\eta_{j}}$
$=125\eta_{j}G^{2}d^{2}D2\tau_{j}+\frac{dD}{\eta_{j}}$
$=126d^{3/2}DG\sqrt{2\tau_{j}}$
where we use $\eta_{j}=1/\sqrt{2dG^{2}\tau_{j}}$. Then, putting everything together and letting $x^{\star}=\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x)$, we get that
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star})$
$\leq\sum_{j=0}^{N-1}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\sum_{j=0}^{N}\sum_{t=\tau_{j}}^{\tau_{j+1}-1}f_{t}(x^{\star}_{j})$
$\leq\sum_{j=0}^{N}126DGd^{3/2}\sqrt{2\tau_{j}}$
$\leq 630DGd^{3/2}\sqrt{2T},$
where we use (9) in the last inequality. ∎
Appendix E Numerical Experiments
In this appendix, we provide further details on the numerical experiments from Section 7 and additional results. Specifically, we discuss the linear-cost setting in Section E.1, the quadratic cost setting in Section E.2.
E.1 Safe Online Linear Programming
First, we give details on the setting in Figure 0(a) and then give additional results.
Algorithm Parameters for Figure 0(a)
First, we give the problem parameters defined by the assumptions:
  • $D=2$
  • $S=\sqrt{2}$
  • $G=\sqrt{2}$
  • $\rho=0.01$
  • $b_{\mathrm{min}}=0.6$
Then, the algorithm parameters for SO-PGD are chosen as suggested by Chaudhary & Kalathil (2022):
  • $\gamma=b_{\mathrm{min}}/S$
  • $\eta=D/(G\sqrt{T})$
  • $T_{0}=T^{2/3}$
  • $\rho=0.01$
  • $\lambda=1$
  • $\delta=0.01$
Also, OSOCO is run with HedgeDescent for $\mathcal{A}$ and the algorithm parameters are chosen as follows:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=0.01$
Note that this choice of $\lambda$ differs from the one specified in Theorem 4.1. However, this choice will only change the regret guarantees by constant factors.
Additional Results
Figure 2: The performance of OGD Zinkevich (2003) with full constraint information and OSOCO in a “hard” setting. The line shows the average regret values over $60$ trials for $T\in\{1000,2000,...,100000\}$ and the shading shows $\pm$ 1 standard deviation.
We also give additional results in a different setting. In particular, we consider a setting that is challenging for online algorithms in general, rather than just safe algorithms. This differs from the settings in Section 7, which are intended to demonstrate the performance of OSOCO in a challenging safe learning setting. The results for OGD Zinkevich (2003) (with full constraint information) and OSOCO are given in Figure 2, where the line shows the regret with $T\in\{1000,2000,...,100000\}$ and is averaged over $60$ trials and the shading shows ±$\pm 1$ standard deviation. The setting is defined as:
  • $d=2$
  • $\mathcal{X}=\sqrt{2}\mathbb{B}$
  • $A=[I\ -I]^{\top}$
  • $\theta_{t}\sim\{-1,1\}^{d}$
  • $\epsilon_{t}\sim\mathcal{U}(0,\rho I)$
  • $\rho=1$
  • $b=\mathbf{1}$
  • $D=2\sqrt{2}$
  • $S=\sqrt{2}$
  • $G=\sqrt{2}$
  • $b_{\mathrm{min}}=1$
Note that this setting is very challenging for both online algorithms in general and safe algorithm in particular. This is evident from the fact that there is a lot of noise on the constraint feedback ($\epsilon_{t}$ has standard deviation $1$), and yet OGD and OSOCO perform similarily.
In this setting, HedgeDescent is used for OSOCO with the following algorithm parameters:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=0.01$
Also, $\eta_{t}=D/(G\sqrt{t})$ is used for OGD Zinkevich (2003).
E.2 Safe Online Quadratic Programming
First, we give the algorithm parameters used for Figure 0(b) and then give some additional results.
Algorithm Parameters for Figure 0(b)
First, we give the problem parameters defined by the assumptions:
  • $D=2$
  • $S=\sqrt{2}$
  • $G=4\sqrt{2}+4$
  • $\rho=0.01$
  • $b_{\mathrm{min}}=0.5$
Then, the algorithm parameters for SO-PGD are chosen as suggested by Chaudhary & Kalathil (2022):
  • $\gamma=b_{\mathrm{min}}/S$
  • $\eta=D/(G\sqrt{T})$
  • $T_{0}=T^{2/3}$
  • $\rho=0.01$
  • $\lambda=1$
  • $\delta=0.01$
Also, OSOCO is run with HedgeDescent for $\mathcal{A}$ and the algorithm parameters are chosen as follows:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=0.01$
Additional Results
Figure 3: The performance of the algorithm in Yu et al. (2017) (abbreviated DPP in reference to “drift-plus penalty”) with full constraint feedback and OSOCO. The points show the average regret values over $30$ trials and the error bars show $\pm$ 1 standard deviation.
Figure 4: The instantaneous violation of the algorithm in Yu et al. (2017) (labeled DPP) with full constraint feedback and OSOCO. The line shows the average regret values over $30$ trials and the shading shows $\pm$ 1 standard deviation.
We also give results comparing OSOCO with the algorithm from Yu et al. (2017). The results of this are shown in Figure 3, where the algorithm from Yu et al. (2017) is referred to as DPP in reference to “drift-plus-penalty”, the name for their analysis technique. The plotted (cumulative) violation is defined as $V_{T}=\sum_{t=1}^{T}(y_{t}-b)$. In this case, we find that the regret of OSOCO is closer to constant $\sqrt{T\log(T)}$ and therefore use that normalization factor. Note that the violation for OSOCO is not shown because it is negative for all trials. We also plot the instantaneous violation for a fixed horizon in Figure 4. The setting is defined as:
  • $f_{t}(x)=3\|x-v_{t}\|^{2}$
  • $d=2$
  • $\mathcal{X}=\mathbb{B}$
  • $A=[-1-1]$
  • $v_{t}\sim[-1,0]^{d}$
  • $\epsilon_{t}\sim\mathcal{N}(0,\rho I)$
  • $\rho=0.01$
  • $b=0.8$
  • $D=2$
  • $S=\sqrt{2}$
  • $G=6\sqrt{2}+6$
  • $b_{\mathrm{min}}=0.8$
In this setting, HedgeDescent is used for OSOCO with the following algorithm parameters:
  • $\eta_{t}=D/(G\sqrt{t})$ (in HedgeDescent)
  • $\zeta_{t}=\sqrt{4\log(2d)}/(GD\sqrt{t})$ (in HedgeDescent)
  • $\lambda=1$
  • $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$
  • $\mu=b_{\mathrm{min}}/T$ (in the sense of Proposition 5.1)
Also, DPP uses the following parameters as recommended in Yu et al. (2017):
  • $\alpha=T$
  • $V=\sqrt{T}$